OpenStudy (anonymous):

please help me....i dislike math lol.... The length of a rectangle is 7 1/2 inches and the width is 2 3/4 inches. What is the ratio, using whole numbers, of the length to the width?

OpenStudy (anonymous):

Length = 7 1/2 inches = 2*7 + 1 / 2 = 15/2 inches and width = 2 3/4 = 2*4 + 3 / 2 = 11/2 inches ratio of Length : Width = 15/2 : 11/2 = 15/2 / 11/2 = 15/2 * 2/11 = 15/11 ratio is 15 : 11

OpenStudy (anonymous):

thankyou